# AtomR: Atomic Operator-Empowered Large Language Models for Heterogeneous Knowledge Reasoning

### v3

@@ 4.1.1. Datasets and Evaluation Metrics
We evaluate AtomR in both single-source and multi-source settings. In the single-source setting, we evaluate AtomR on three Wikipedia-based multi-hop QA benchmarks: HotpotQA (Yang et al., 2018), 2WikiMultiHop (Ho et al., 2020), and Musique (Trivedi et al., 2022). We adopt the test and development sets released by IRCoT (Trivedi et al., 2023), which include a 500-entry test set and a 100-entry development set. In the multi-source setting, we evaluate AtomR on two recent datasets: CRAG (Yang et al., 2024) by Meta and BlendQA by us. Both CRAG and BlendQA are constructed on three heterogeneous knowledge sources—the Wikidata knowledge base, the Wikipedia text corpus, and the Google Web search engine. We will provide a more thorough introduction of BlendQA in the next section. For BlendQA, we evaluate AtomR on the full 445-entry dataset; for CRAG, we only consider static questions and sample a 500-entry test set with a 100-entry development set. Following IRCoT and ProbTree, we adopt token-level F1 as the evaluation metric.
@@ 4.1.2. BlendQA
We propose BlendQA, a dataset to evaluate models’ cross-knowledge source reasoning capabilities. Specifically, we construct BlendQA on three heterogeneous knowledge sources: a full Wikipedia dump as the text source, Google as the web source, and Wikidata as the knowledge graph. We use gpt-4o-2024-08-06 as the LLM to aid dataset construction. The general construction process is to generate two sub-questions $sub$-$q_{1}$ and $sub$-$q_{2}$ from two different sources that share a common bridging entity, and then merge them together to form a cohesive question. There are three types of questions: KG-Text, KG-Web, Text-Web.
KG-Text. For KG-Text questions, $sub$-$q_{1}$ is sampled from NaturalQuestions (NQ) (Kwiatkowski et al., 2019), which is built on Wikipedia. We first identify $sub$-$q_{1}$’s topic entity as the bridging entity, then sample the topic entity’s triples from the KG to generate $sub$-$q_{2}$.
KG-Web. There are two sub-types of KG-Web questions depending on the anchor of the constructed data. (1) KG2Web is anchored in the KG. Specifically, we randomly sample entities from the KG as bridging entities, and sample their triples to generate $sub$-$q_{1}$. Then, we search the bridging entity for relevant news to generate $sub$-$q_{2}$. (2) Web2KG is anchored in the web. Specifically, we collect a large number of news articles from various domains (politics, business, etc.), and ask the LLM to identify a bridging entity in each article and ask a question about it ($sub$-$q_{1}$). Then, we generate $sub$-$q_{2}$ by sampling triples from the KG using the bridging entity.
Text-Web. There are also two sub-types of Text-Web questions. (1) Web2Text. We sample $sub$-$q_{1}$ from NQ and identify the topic entity as the bridging entity. Then, we search the entity on the web (excluding Wikipedia articles) to collect its descriptions. With these descriptions, we use the LLM to generate a unique tag of the entity to serve as $sub$-$q_{2}$, e.g. “Neil Armstrong” - “first man to walk on the moon”. (2) Text2Web. Again, we sample $sub$-$q_{1}$ from NQ, and use the answer of it as the bridging entity (if valid). Then we search the entity’s news articles to generate $sub$-$q_{2}$.
@@ 4.1.4. Implementation Details
For all experiments, we use gpt-4o-2024-08-06 as our base LLM. For web knowledge retrieval, we use SERPAPI to access Google. For text knowledge retrieval, we follow previous works (Yoran et al., 2023; Xu et al., 2024) to use ColBERTv2 as our text retriever. We use HotpotQA’s official Wikipedia abstract dump from October 2017 to evaluate HotpotQA, and the Dec 2021 full Wikipedia dump from Atlas (Izacard et al., 2023) to evaluate the other four datasets. For KB knowledge retrieval, we use KoPL as the query language and KQA Pro (Cao et al., 2022)’s Wikidata dump as the knowledge base. We only retrieve text knowledge for single-source datasets, and retrieve knowledge from as many sources as supported for each baseline on multi-source datasets. We retrieve knowledge from all three sources for multi-source Standard RAG. For AtomR, we set $k=3$ to retrieve the top 3 web and text results, and set $t=0.5$ for the Filter function. The results for IRCoT are from Cao et al. (Cao et al., 2023), while all other baselines are reproduced by ourselves using the above settings.
@@ 4.3.1. Comparative Case Study: How does AtomR outperform previous methods?
Figure 3. Comparative case study of AtomR with previous SOTA Chain-of-Knowledge and ProbTree.
[…]
@@ 4.3.2. Effect of Incorporating Multi-Source Knowledge
w/o KG 46.52$(\downarrow 0.5)$ 68.88$(\downarrow 1.4)$
w/o Text 43.34$(\downarrow 3.6)$ 67.03$(\downarrow 3.2)$
w/o Web 30.30$(\downarrow 16.7)$ 58.27$(\downarrow 12.0)$
@@ 4.3.3. Comparative Cost Analysis
Figure 4. Comparing AtomR’s consumption costs with the previous SOTA ProbTree.
[…]
@@ Appendix A Algorithm for Atomic Reasoning Execution
1:Require: Atomic Reasoning Tree $ART$
2:for $n_{i}$ in PostOrder($ART$do
@@ Appendix B Distribution of Atomic Knowledge Operators
Figure 5 illustrates the test-time distribution of the three atomic knowledge operators for AtomR.
Figure 5. Distribution of atomic knowledge operator calls.
[…]
We observe that Relate is the most frequently selected operator across all datasets, which is expected, as it serves for one-hop inference—the fundamental operation for addressing multi-hop questions. In contrast, Filter is the least selected operator, with particularly low selection in Musique, where it is chosen only $0.6\%$ of the time. However, this is notable because Musique’s data construction process did not involve entity filtering operations. The minimal selection of Filter on Musique actually demonstrates AtomR’s ability to intelligently adapt to differently constructed datasets by autonomously adjusting the use of the three operators.
@@ Appendix C Distribution of Selected Knowledge Sources
Figure 6 displays the test-time distribution of the three heterogeneous knowledge sources on the two muti-source datasets.
Figure 6. Distribution of selected knowledge sources.
[…]
This distribution aligns highly to the actual structure of the datasets: BlendQA is constructed on balanced data from all three sources, while CRAG is constructed mainly on text and web sources. This demonstrates that AtomR yields a promising knowledge source selection strategy.
@@ Appendix D Ablating backbone LLMs
AtomR 72.06$(\uparrow 14.7)$ 79.71$(\uparrow 29.2)$ 33.47$(\uparrow 26.9)$
@@
AtomR 56.72$(\uparrow 7.3)$ 72.15$(\uparrow 29.7)$ 29.92$(\uparrow 15.5)$
@@
AtomR 23.99$(\downarrow 24.0)$ 20.77$(\downarrow 10.7)$ 6.32$(\downarrow 2.9)$
Table 5 illustrates the performance of AtomR across three backbone LLMs: GPT-4o (the LLM choice of our main experiments), Llama-3.1-70B-Instruct, and Llama-3.1-8B-Instruct. Results show that AtomR yields outstanding performance gains with Llama-3.1-70b-Instruct as the backbone, achieving significant F1 gains over the Standard RAG baseline on all three datasets (as high as $29.7\%$ on 2WikiMultiHop). However, AtomR leads to serious performance degradation with Llama-3.1-8b-Instruct. This shows that AtomR is a framework too complex for LLMs of small sizes. Yet, AtomR’s success with Llama-3.1-70b-Instruct demonstrates that it could yield promising results with 70b-level open-source models, which are widely accessible to the research community.
@@ Appendix E Ablating text retriever models
AtomR is a plug-and-play framework that works with any text retriever model. To investigate the effect of text retriever choice on AtomR’s performance, we also conduct an ablation on AtomR’s text retriever model. We use the same 100-entry subsets from Appendix D.
@@
Table 6 illustrates the performance of AtomR using three text retriever models: ColBERTv2 (the retriever choice of our main experiments), BGE-large-en-v1.5, and BM25. Reslts show that AtomR yields comparative performance among the two dense retriever models ColBERTv2 and BGE-v1.5, but shows a slight performance degradation on the term-based retrieval model BM25. Thus, it is best to implement AtomR with dense text retriever models to optimize its performance.
Appendix F In-context Learning Prompts
Figures  7 to  13 display all LLM prompts used for AtomR.
Figure 7. Prompt for the adaptive LLM executor of the Search operator.
[…]
Figure 8. Prompt for the adaptive LLM executor of the Relate operator.
[…]
Figure 9. Prompt for the adaptive LLM executor of the Filter operator.
[…]
Figure 10. Prompt for Atomic Reasoning Tree generation.
[…]
Figure 11. Prompt for Dynamic Knowledge Source Selection.
[…]
Figure 12. Prompt for Child and Sibling Answer Reasoning.
[…]
Figure 13. Prompt for Direct RAG Reasoning.
[…]

### v4

@@ 4.1.1. Datasets and Evaluation Metrics
We evaluate AtomR in both single-source and multi-source settings. In the single-source setting, we evaluate AtomR on three Wikipedia-based multi-hop QA benchmarks: HotpotQA (Yang et al., 2018), 2WikiMultiHop (Ho et al., 2020), and Musique (Trivedi et al., 2022). We adopt the test and development sets released by IRCoT (Trivedi et al., 2023), which include a 500-entry test set and a 100-entry development set. In the multi-source setting, we evaluate AtomR on two recent datasets: CRAG (Yang et al., 2024b) by Meta and BlendQA by us. Both CRAG and BlendQA are constructed on three heterogeneous knowledge sources—the Wikidata knowledge base, the Wikipedia text corpus, and the Google Web search engine. We will provide a more thorough introduction of BlendQA in the next section. For BlendQA, we evaluate AtomR on the full 445-entry dataset; for CRAG, we only consider static questions and sample a 500-entry test set with a 100-entry development set. Following IRCoT and ProbTree, we adopt token-level F1 as the evaluation metric.
@@ 4.1.2. BlendQA
We propose BlendQA, a dataset to evaluate models’ cross-knowledge source reasoning capabilities. Specifically, we construct BlendQA on three heterogeneous knowledge sources: a full Wikipedia dump as the text source, Google as the web source, and Wikidata as the knowledge graph. We use gpt-4o-2024-08-06 as the LLM to aid dataset construction. The general construction process is to generate two sub-questions subsub-$q_{1}$ and subsub-$q_{2}$ from two different sources that share a common bridging entity, and then merge them together to form a cohesive question. There are three types of questions: KG-Text, KG-Web, Text-Web.
KG-Text. For KG-Text questions, subsub-$q_{1}$ is sampled from NaturalQuestions (NQ) (Kwiatkowski et al., 2019), which is built on Wikipedia. We first identify subsub-$q_{1}$’s topic entity as the bridging entity, then sample the topic entity’s triples from the KG to generate subsub-$q_{2}$.
KG-Web. There are two sub-types of KG-Web questions depending on the anchor of the constructed data. (1) KG2Web is anchored in the KG. Specifically, we randomly sample entities from the KG as bridging entities, and sample their triples to generate subsub-$q_{1}$. Then, we search the bridging entity for relevant news to generate subsub-$q_{2}$. (2) Web2KG is anchored in the web. Specifically, we collect a large number of news articles from various domains (politics, business, etc.), and ask the LLM to identify a bridging entity in each article and ask a question about it (subsub-$q_{1}$). Then, we generate subsub-$q_{2}$ by sampling triples from the KG using the bridging entity.
Text-Web. There are also two sub-types of Text-Web questions. (1) Web2Text. We sample subsub-$q_{1}$ from NQ and identify the topic entity as the bridging entity. Then, we search the entity on the web (excluding Wikipedia articles) to collect its descriptions. With these descriptions, we use the LLM to generate a unique tag of the entity to serve as subsub-$q_{2}$, e.g. “Neil Armstrong” - “first man to walk on the moon”. (2) Text2Web. Again, we sample subsub-$q_{1}$ from NQ, and use the answer of it as the bridging entity (if valid). Then we search the entity’s news articles to generate subsub-$q_{2}$.
@@ 4.1.4. Implementation Details
For all experiments, we use gpt-4o-2024-08-06 as our base LLM. For web knowledge retrieval, we use SERPAPI to access Google. For text knowledge retrieval, we follow previous works (Yoran et al., 2023; Xu et al., 2024) to use ColBERTv2 as our text retriever. We use HotpotQA’s official Wikipedia abstract dump from October 2017 to evaluate HotpotQA, and the Dec 2021 full Wikipedia dump from Atlas (Izacard et al., 2023b) to evaluate the other four datasets. For KB knowledge retrieval, we use KoPL as the query language and KQA Pro (Cao et al., 2022)’s Wikidata dump as the knowledge base. We only retrieve text knowledge for single-source datasets, and retrieve knowledge from as many sources as supported for each baseline on multi-source datasets. We retrieve knowledge from all three sources for multi-source Standard RAG. For AtomR, we set $k=3$ to retrieve the top 3 web and text results, and set $t=0.5$ for the Filter function. The results for IRCoT are from Cao et al. (Cao et al., 2023), while all other baselines are reproduced by ourselves using the above settings.
@@ 4.3.1. Comparative Case Study: How does AtomR outperform previous methods?
Figure 3. Comparative case study of AtomR with previous SOTA Chain-of-Knowledge and ProbTree.:
@@ 4.3.2. Effect of Incorporating Multi-Source Knowledge
w/o KG 46.52(0.5)($\downarrow$ 0.5) 68.88(1.4)($\downarrow$ 1.4)
w/o Text 43.34(3.6)($\downarrow$ 3.6) 67.03(3.2)($\downarrow$ 3.2)
w/o Web 30.30(16.7)($\downarrow$ 16.7) 58.27(12.0)($\downarrow$ 12.0)
@@ 4.3.3. Comparative Cost Analysis
Figure 4. Comparing AtomR’s consumption costs with the previous SOTA ProbTree.:
@@ 5. Conclusion
Acknowledgements.
This work is supported by National Natural Science Foundation of China (62476150), Beijing Natural Science Foundation (L243006), and Tsinghua University Initiative Scientific Research Program.
@@ Appendix A Algorithm for Atomic Reasoning Execution
1:Require: Atomic Reasoning Tree ARTART
2:for $n_{i}$ in PostOrder(ARTART) do
@@ Appendix B Distribution of Atomic Knowledge Operators
Figure 5 illustrates the test-time distribution of AtomR’s three atomic knowledge operators. Relate is the most frequently selected operator across all datasets, which is expected, as it serves for one-hop inference—the fundamental operation of multi-hop reasoning. In contrast, Filter is the least selected operator, with particularly low selection in Musique, where it is chosen only $0.6\%$ of the time. However, this is notable because Musique’s data construction process did not involve entity filtering operations. The minimal selection of Filter on Musique demonstrates AtomR’s ability to intelligently adapt to differently constructed datasets by autonomously adjusting the use of the three operators.
Figure 5. Distribution of atomic knowledge operator calls.:
@@ Appendix C Distribution of Selected Knowledge Sources
Figure 6 displays the test-time distribution of the three heterogeneous knowledge sources on the two muti-source datasets. This distribution aligns highly to the actual structure of the datasets: BlendQA is constructed on balanced data from all three sources, while CRAG is constructed mainly on text and web sources. This demonstrates that AtomR yields a promising knowledge source selection strategy.
@@ Appendix D Ablating backbone LLMs
AtomR 72.06(14.7)($\uparrow$ 14.7) 79.71(29.2)($\uparrow$ 29.2) 33.47(26.9)($\uparrow$ 26.9)
@@
AtomR 56.72(7.3)($\uparrow$ 7.3) 72.15(29.7)($\uparrow$ 29.7) 29.92(15.5)($\uparrow$ 15.5)
@@
AtomR 23.99(24.0)($\downarrow$ 24.0) 20.77(10.7)($\downarrow$ 10.7) 6.32(2.9)($\downarrow$ 2.9)
Table 5 illustrates the performance of AtomR across three backbone LLMs: GPT-4o, Llama-3.1-70B-Instruct, and Llama-3.1-8B-Instruct. Results show that AtomR yields outstanding performance gains with Llama-3.1-70b-Instruct as the backbone, achieving significant F1 gains over the Standard RAG baseline on all three datasets. However, AtomR leads to serious performance degradation with Llama-3.1-8b-Instruct. This shows that AtomR is a framework too complex for LLMs of small sizes. Yet, AtomR’s success with Llama-3.1-70b-Instruct demonstrates that it could yield promising results with 70b-level open-source models, which are widely accessible to the research community.
Figure 6. Distribution of selected knowledge sources.:
@@ Appendix E Ablating text retriever models
AtomR is a plug-and-play framework that works with any text retriever model. We conduct an ablation on AtomR’s text retriever model using the same 100-entry subsets from Appendix D. Table 6 illustrates the performance of AtomR using three text retriever models: ColBERTv2, BGE-large-en-v1.5, and BM25.
@@
Results show that AtomR yields comparative performance among the two dense retriever models ColBERTv2 and BGE-v1.5, but shows a slight performance degradation on the term-based retrieval model BM25. Thus, it is best to implement AtomR with dense text retriever models to optimize its performance.
Appendix F Comparing AtomR to Additional Baselines
Considering that AtomR is a RAG framework that supports (1) dynamic retrieval, and (2) KG-level atomic knowledge manipulation, we compare it against two additional baselines: (1) DRAGIN(Su et al., 2024), a dynamic RAG framework with real-time retrieval decisions; (2) HippoRAG(Gutiérrez et al., 2024), a RAG framework that combines LLM KG construction and KG-based RAG. We experiment on the 100-entry subsets from Appendix D with Llama-3.1-70b-Instruct as the backbone.
Table 7. Comparing AtomR to additional baselines
HotpotQA 2Wiki Musique Avg.
DRAGIN (Su et al., 2024) 23.4 56.6 19.3 33.1
HippoRAG (Gutiérrez et al., 2024) 18.7 14.4 30.3 21.1
AtomR 56.7 72.2 29.9 52.9
Table 7 reports the F1 scores of DRAGIN, HippoRAG, and AtomR. Overall, AtomR significantly outperforms both methods, achieving respective F1 gains of 31.8% and 19.8%. Furthermore, on the HotpotQA subset, AtomR outperforms HippoRAG by achieving 303.2% F1 performance with only 2.5% completion token consumption. This is because AtomR achieves KG-level fine-grained knowledge retrieval and reasoning without HippoRAG’s expensive construction of an actual KG. These results highlight the effectiveness and cost-efficiency of AtomR ’s framework and operator design.
Appendix G AtomR Tree Depth Analysis
To better understand the computational complexity of AtomR, we conduct a tree depth analysis of AtomR’s Atomic Reasoning Trees (ARTs). As shown in Table 8, AtomR produces ARTs with controllable complexity, with a maximum depth of 5 on HotpotQA and an average depth below 3 across all datasets. Notably, the CRAG dataset exhibits the shallowest tree depth distribution, reflecting the simplicity of its questions compared to the other datasets. These results indicate that AtomR is capable of adapting its tree depth dynamically to match the complexity of different questions.
Table 8. AtomR tree depth analysis.
Dataset Mean Median Min Max
HotpotQA 2.65 3.00 2.00 5.00
2Wiki 2.97 3.00 2.00 3.00
Musique 2.47 2.00 2.00 4.00
BlendQA 2.42 2.00 2.00 4.00
CRAG 1.84 2.00 1.00 3.00
Figure 7. Prompt for Atomic Reasoning Tree generation. :
Appendix H In-context Learning Prompts
Figures 7 to 9 display representative LLM prompts of AtomR. For all prompts, please refer to our official code repository.
Figure 8. Prompt for the adaptive LLM executor of the Search operator.:
Figure 9. Prompt for Dynamic Knowledge Source Selection.: